OpenStudy (anonymous):

List the potential rational zeros of the following function. Please show all work b/c I am not understanding this. f(x)=7x^4-x^2+3

OpenStudy (mertsj):

The possible rational roots are the factors of the constant term (which in this case is 3) divided by the factors of the coefficient of the leading term (which in this case is 7). So the possible rational roots are: \[\frac{\pm1,\pm3}{\pm1,\pm7}\]

OpenStudy (mertsj):

You can then use synthetic division to see if any of them are actual roots.

OpenStudy (anonymous):

what do u mean

OpenStudy (mertsj):

do you know synthetic division?

OpenStudy (mertsj):

I can do an example. Perhaps it will jog your memory.

OpenStudy (mertsj):

Oh. Perhaps you have not yet learned it because the problem says to list the possible rational zeros. They are as I have indicated. Do you want a complete list?

OpenStudy (mertsj):

\[\pm1,\pm3,\pm\frac{1}{7},\pm\frac{3}{7}\]

OpenStudy (anonymous):

so do i divide them, put them into an equation or what?

OpenStudy (mertsj):

I think you are done. The problem only asked you for the possibilities.

OpenStudy (anonymous):

oh ok

OpenStudy (mertsj):

I think after you master this your next lesson might be to see whether any of the possibilities are actual roots. Then you will probably learn synthetic division and some other techniques for finding roots of polynomial equations.